# Quantum backflow in biased tight-binding systems

### v1

@@ Appendix A Derivation of optimal bounds on the flux for the periodic chain
First, we start defining the functional from the expression of the flux at a fixed $j^{\prime},t^{\prime}$ and the constraint defined by the normalization condition eq.(51)
$\scriptstyle I_{1}(c_{m}^{*},c_{n})=\frac{2\tau\sqrt{1+\epsilon^{2}}}{N\hbar}\sum_{n=\eta_{1}}^{\eta_{2}}\sum_{m=\eta_{1}}^{\eta_{2}}c^{*}_{m}c_{n}\,e^{i\left(\frac{(2j^{\prime}-1)(m-n)\pi}{N}-\frac{(E_{n}-E_{m})t^{\prime}}{\hbar}\right)}\sin\left(\frac{(m+n)\pi}{N}+\xi\right)-\lambda\sum_{m=\eta_{1}}^{\eta_{2}}|c_{m}|^{2}$ (68)
@@
$\frac{2\tau\sqrt{1+\epsilon^{2}}}{N\hbar}\sum_{n=\eta_{1}}^{\eta_{2}}c_{n}\,e^{i\left(\frac{(2j^{\prime}-1)(m-n)\pi}{N}-\frac{(E_{n}-E_{m})t^{\prime}}{\hbar}\right)}\sin\left(\frac{(m+n)\pi}{N}+\xi\right)=\lambda c_{m}$ (69)
@@
$\begin{split}\scriptstyle\frac{2\tau\sqrt{1+\epsilon^{2}}}{N\hbar}e^{-i\left(\frac{(2j^{\prime}-1)m\pi}{N}-\frac{E_{m})t^{\prime}}{\hbar}\right)}\Big[\cos\left(\frac{m}{N}\pi+\frac{\xi}{2}\right)\sum_{n=\eta_{1}}^{\eta_{2}}c_{n}\,e^{i\left(\frac{(2j^{\prime}-1)n\pi}{N}-\frac{E_{n}t^{\prime}}{\hbar}\right)}\sin\left(\frac{n}{N}\pi+\frac{\xi}{2}\right)\\ \scriptstyle+\sin\left(\frac{m}{N}\pi+\frac{\xi}{2}\right)\sum_{n=\eta_{1}}^{\eta_{2}}c_{n}\,e^{i\left(\frac{(2j^{\prime}-1)n\pi}{N}-\frac{E_{n}t^{\prime}}{\hbar}\right)}\cos\left(\frac{n}{N}\pi+\frac{\xi}{2}\right)\Big]=\lambda c_{m}\end{split}$ (70)
To solve the last equation, we set
$\begin{split}\sum_{n=\eta_{1}}^{\eta_{2}}c_{n}\,e^{i\left(\frac{(2j^{\prime}-1)n\pi}{N}-\frac{E_{n}t^{\prime}}{\hbar}\right)}\sin\left(\frac{n}{N}\pi+\frac{\xi}{2}\right)=\alpha\\ \sum_{n=\eta_{1}}^{\eta_{2}}c_{n}\,e^{i\left(\frac{(2j^{\prime}-1)n\pi}{N}-\frac{E_{n}t^{\prime}}{\hbar}\right)}\cos\left(\frac{n}{N}\pi+\frac{\xi}{2}\right)=\beta\end{split}$ (71)
@@
$c_{m}=\frac{2\tau\sqrt{1+\epsilon^{2}}}{N\hbar\lambda}e^{-i\left(\frac{(2j^{\prime}-1)m\pi}{N}-\frac{E_{m})t^{\prime}}{\hbar}\right)}\Big[\alpha\cos\left(\frac{m}{N}\pi+\frac{\xi}{2}\right)+\beta\sin\left(\frac{m}{N}\pi+\frac{\xi}{2}\right)\Big].$ (72)
@@
$\begin{split}\frac{2\tau\sqrt{1+\epsilon^{2}}}{N\hbar\lambda}\sum_{n=\eta_{1}}^{\eta_{2}}\Big[\alpha\cos\left(\frac{n}{N}\pi+\frac{\xi}{2}\right)+\beta\sin\left(\frac{n}{N}\pi+\frac{\xi}{2}\right)\Big]\sin\left(\frac{n}{N}\pi+\frac{\xi}{2}\right)=\alpha\\ \frac{2\tau\sqrt{1+\epsilon^{2}}}{N\hbar\lambda}\sum_{n=\eta_{1}}^{\eta_{2}}\Big[\alpha\cos\left(\frac{n}{N}\pi+\frac{\xi}{2}\right)+\beta\sin\left(\frac{n}{N}\pi+\frac{\xi}{2}\right)\Big]\cos\left(\frac{n}{N}\pi+\frac{\xi}{2}\right)=\beta\end{split}$ (73)
@@
$\mathbf{A}\begin{pmatrix}\alpha\\ \beta\end{pmatrix}=\lambda\begin{pmatrix}\alpha\\ \beta\end{pmatrix}$ (74)
@@
$\beta=\sqrt{\frac{\eta_{2}+1-\eta_{1}+\cos\!\left(\tfrac{\pi(\eta_{1}+\eta_{2})}{N}+\xi\right)\csc\!\left(\tfrac{\pi}{N}\right)\sin\!\left(\tfrac{\pi(\eta_{2}+1-\eta_{1})}{N}\right)}{\mathcal{D_{\pm}}}}$ (80)

### v2

@@ Appendix A Derivation of optimal bounds on the flux for the periodic chain
First, we define the functional from the expression for the flux at a fixed $j^{\prime},t^{\prime}$ and the constraint defined by the normalization condition eq.(51)
$\scriptstyle I_{1}(c_{m}^{*},c_{n})=\frac{2\tau\sqrt{1+\epsilon^{2}}}{N\hbar}\sum_{n=\eta_{1}}^{\eta_{2}}\sum_{m=\eta_{1}}^{\eta_{2}}c^{*}_{m}c_{n}\,e^{i\left(\frac{(2j^{\prime}-1)(m-n)\pi}{N}-\frac{(E_{n}-E_{m})t^{\prime}}{\hbar}\right)}\sin\left(\frac{(m+n)\pi}{N}+\xi\right)-\lambda\sum_{m=\eta_{1}}^{\eta_{2}}|c_{m}|^{2}.$ (70)
@@
$\frac{2\tau\sqrt{1+\epsilon^{2}}}{N\hbar}\sum_{n=\eta_{1}}^{\eta_{2}}c_{n}\,e^{i\left(\frac{(2j^{\prime}-1)(m-n)\pi}{N}-\frac{(E_{n}-E_{m})t^{\prime}}{\hbar}\right)}\sin\left(\frac{(m+n)\pi}{N}+\xi\right)=\lambda c_{m},$ (71)
@@
$\begin{split}\scriptstyle\frac{2\tau\sqrt{1+\epsilon^{2}}}{N\hbar}e^{-i\left(\frac{(2j^{\prime}-1)m\pi}{N}-\frac{E_{m})t^{\prime}}{\hbar}\right)}\Big[\cos\left(\frac{m}{N}\pi+\frac{\xi}{2}\right)\sum_{n=\eta_{1}}^{\eta_{2}}c_{n}\,e^{i\left(\frac{(2j^{\prime}-1)n\pi}{N}-\frac{E_{n}t^{\prime}}{\hbar}\right)}\sin\left(\frac{n}{N}\pi+\frac{\xi}{2}\right)\\ \scriptstyle+\sin\left(\frac{m}{N}\pi+\frac{\xi}{2}\right)\sum_{n=\eta_{1}}^{\eta_{2}}c_{n}\,e^{i\left(\frac{(2j^{\prime}-1)n\pi}{N}-\frac{E_{n}t^{\prime}}{\hbar}\right)}\cos\left(\frac{n}{N}\pi+\frac{\xi}{2}\right)\Big]=\lambda c_{m}.\end{split}$ (72)
To solve this equation, we set
$\begin{split}\sum_{n=\eta_{1}}^{\eta_{2}}c_{n}\,e^{i\left(\frac{(2j^{\prime}-1)n\pi}{N}-\frac{E_{n}t^{\prime}}{\hbar}\right)}\sin\left(\frac{n}{N}\pi+\frac{\xi}{2}\right)=\alpha,\\ \sum_{n=\eta_{1}}^{\eta_{2}}c_{n}\,e^{i\left(\frac{(2j^{\prime}-1)n\pi}{N}-\frac{E_{n}t^{\prime}}{\hbar}\right)}\cos\left(\frac{n}{N}\pi+\frac{\xi}{2}\right)=\beta,\end{split}$ (73)
@@
$c_{m}=\frac{2\tau\sqrt{1+\epsilon^{2}}}{N\hbar\lambda}e^{-i\left(\frac{(2j^{\prime}-1)m\pi}{N}-\frac{E_{m}t^{\prime}}{\hbar}\right)}\Big[\alpha\cos\left(\frac{m}{N}\pi+\frac{\xi}{2}\right)+\beta\sin\left(\frac{m}{N}\pi+\frac{\xi}{2}\right)\Big].$ (74)
@@
$\begin{split}\frac{2\tau\sqrt{1+\epsilon^{2}}}{N\hbar\lambda}\sum_{n=\eta_{1}}^{\eta_{2}}\Big[\alpha\cos\left(\frac{n}{N}\pi+\frac{\xi}{2}\right)+\beta\sin\left(\frac{n}{N}\pi+\frac{\xi}{2}\right)\Big]\sin\left(\frac{n}{N}\pi+\frac{\xi}{2}\right)=\alpha,\\ \frac{2\tau\sqrt{1+\epsilon^{2}}}{N\hbar\lambda}\sum_{n=\eta_{1}}^{\eta_{2}}\Big[\alpha\cos\left(\frac{n}{N}\pi+\frac{\xi}{2}\right)+\beta\sin\left(\frac{n}{N}\pi+\frac{\xi}{2}\right)\Big]\cos\left(\frac{n}{N}\pi+\frac{\xi}{2}\right)=\beta,\end{split}$ (75)
@@
$\mathbf{A}\begin{pmatrix}\alpha\\ \beta\end{pmatrix}=\lambda\begin{pmatrix}\alpha\\ \beta\end{pmatrix}$, (76)
@@
$\beta=\sqrt{\frac{\eta_{2}+1-\eta_{1}+\cos\!\left(\tfrac{\pi(\eta_{1}+\eta_{2})}{N}+\xi\right)\csc\!\left(\tfrac{\pi}{N}\right)\sin\!\left(\tfrac{\pi(\eta_{2}+1-\eta_{1})}{N}\right)}{\mathcal{D_{\pm}}}},$ (82)